How many ounces of a 25% alcohol solution must be mixed with a 11 ounces of a 30% of a solution to make a 26% alcohol solution
Answer by richwmiller(17219)

We have 11 oz of 30% alcohol
We want to add some 25% alcohol to make 26% alcohol
0.25*x+0.3*11=0.26(11+x)
0.25*x+3.3=2.86+0.26x
0.25x-0.26x=2.86-3.3
-0.01x=-0.44
x=44 oz at 25% (11) is added to 11 oz at 30%(3.3) to make 26% of 55 oz(14.3)
check
0.25*44+0.3*11=0.26(11+44)
11+3.3=0.26(55)
14.3=14.3
ok